OpenStudy (anonymous):

If a company deposits money into an account today that pays 6% annual interest compounded quarterly. Find the amount invested to total 55,000 in 5 years

OpenStudy (kropot72):

The quarterly interest is 6/4 = 1.5% The number of quarters in 5 years is 5 * 4 = 20 \[55000=P(1+0.015)^{20}\] CAn u now solve for P the amount invested?

OpenStudy (anonymous):

40835.87

OpenStudy (anonymous):

am I right?

OpenStudy (kropot72):

Congratulations, u solved it :)

OpenStudy (ledah):

http://web2.0calc.com/ use this calculator

OpenStudy (ledah):

for future math equations